# Perch 2.0: The Bittern Lesson for Bioacoustics

### v1

@@ 1 Introduction
In this paper we introduce Perch 2.0, a new iteration of the Perch model that achieves state-of-the-art results on two bioacoustics benchmarks: BirdSET and the BEnchmark of Animal Sounds (BEANS) (Section 3). Perch 2.0 was primarily trained with the task of species classification. Although self-supervised learning has been explored in the literature as a way to further improve bioacoustics models—Bird-MAE (Rauch et al., 2025), BirdAVES (Hagiwara, 2023) and SimCLR-style models (Moummad et al., 2024)—these approaches generally struggle to outperform strong supervised baselines (Ghani et al., 2023; Kather et al., 2025). In our exploratory research we too have experimented with a variety of self-supervised methods such as MAEs, HuBERT and SimCLR but experienced a similar inability to consistently outperform supervised models. We present some hypotheses as to why the fine-grained task of (avian) species classification is a particularly useful pre-training task (Section 4).
Compared to previous iterations of the Perch model we use additional training data (including non-avian taxa; Section 2.1) and new data augmentations and training objectives. We generally find that increasing the difficulty of the classification problem increases the overall quality of the embedding model: To that end, we introduce a novel generalization of mixup (Zhang et al., 2018) that mixes more than two sources (Section 2.1). Additionally, recent BirdCLEF competitions have demonstrated strong results with iterative pseudo-labeling (Kahl et al., 2024). Building on this observation, we introduce a variation of self-distillation (Allen-Zhu and Li, 2022) where a prototype learning classifier (Chen et al., 2019; Heinrich et al., 2025) is used to generate predictions that are used as soft targets for the network’s dense linear classifier (Section 2.2 and Section 2.3). Finally, noting that label granularity contributes to transfer learning performance (Section C.1), we add an auxiliary self-supervised loss in the form of source prediction (DIET, Balestriero, 2023), asking the model to predict the source recording of an audio window (Section 2.2 and Section 2.3).
@@ 2.5 Evaluation
The field of bioacoustics recognized domain shift as an important problem early on (Lasseck and Others, 2013), acknowledging that deployment conditions and tasks might differ from training time. Hence, our model selection (validation) is set up to test several forms of generalization: we consider avian soundscapes (which are qualitatively different from the focal recordings found in our training data); consider tasks other than species identification (e.g., call-type and dialect recognition); and evaluate transfer to species classification of non-avian taxa (bats, marine mammals, mosquitoes, etc.) For our model evaluation (testing) we use both BirdSet (Rauch et al., 2024) and BEANS (Hagiwara et al., 2023). Combined, these benchmarks cover a similar set of domain shifts while still allowing us to make a fair comparison with existing models.
In practice, bioacoustics models like Perch are often used by practitioners who have limited computational resources and machine learning expertise while working on novel problems with little or no labeled data (but potentially large amounts of unlabeled data). In these settings a sensible approach is to embed the data once and then use methods such as clustering, nearest neighbor searches, few-shot learning and agile learning (Dumoulin et al., 2025). We want to ensure that our model produces embeddings that perform well in these situations, so during both model selection and evaluation we freeze the embedding model.
@@ 2.5.1 Model-selection tasks
    We evaluate pretrained classifier performance using ROC-AUC on two fully-annotated bird datasets, where the class labels are a subset of the training classes of the Perch model: Powdermill (Denton et al., 2022) (as in Rauch et al. (2024)) and Caples (Denton et al., 2022). These tasks validate that our model makes useful species predictions out-of-the-box. We make predictions for every 5 s window with a 2.5 s stride and count the prediction as correct when it overlaps with a ground truth annotation.
@@
    Finally, we have linear transfer tasks in which we select a random subset of $16$ examples per class and compute an embedding for each (by embedding each 5 s window with a stride of 5 s and averaging them). We then train a linear classifier for 10,000 steps on the selected examples using scikit-learn. All remaining examples are used to compute an ROC-AUC score. This follows the procedure applied in Ghani et al. (2023) and evaluates how well our model’s embeddings do in few-shot and agile learning setups. We evaluate this task on the BEANS classification tasks, some of the datasets from Ghani et al. (2023) as well as marine datasets such as DCLDE (Palmer et al., 2025), NOAA (NOAA Pacific Islands Fisheries Science Center, 2021) and ReefSet (Williams et al., 2025).
@@
Powdermill Chronister et al. (2021); Rauch et al. (2024)
@@
BirdSet Rauch et al. (2024); Alexander Hopping et al. (2022); Vega-Hidalgo et al. (2023); Kahl et al. (2022a); Clapp et al. (2023); Kahl et al. (2022b)
@@ 2.5.2 Evaluation tasks
The BirdSet benchmark (Rauch et al., 2024) consists of six fully-annotated datasets from the continental United States, Hawai‘i, Peru and Columbia. BirdSet provides training sets for each of these soundscapes for model fine-tuning. In our evaluations we forgo fine-tuning and simply use the predictions of our prototype learning classifier predictions directly.

### v2

@@ 1 Introduction
In this paper we introduce Perch 2.0, a new iteration of the Perch model that achieves state-of-the-art results on two bioacoustics benchmarks: BirdSET and the BEnchmark of ANimal Sounds (BEANS) (Section 3). Perch 2.0 was primarily trained with the task of species classification. Although self-supervised learning has been explored in the literature as a way to further improve bioacoustics models—Bird-MAE (Rauch et al., 2025a), BirdAVES (Hagiwara, 2023) and SimCLR-style models (Moummad et al., 2024)—these approaches generally struggle to outperform strong supervised baselines (Ghani et al., 2023; Kather et al., 2025). In our exploratory research we too have experimented with a variety of self-supervised methods such as MAEs, HuBERT and SimCLR but experienced a similar inability to consistently outperform supervised models. We present some hypotheses as to why the fine-grained task of (avian) species classification is a particularly useful pre-training task (Section 4).
Compared to previous iterations of the Perch model we use additional training data (including non-avian taxa; Section 2.1) and new data augmentations and training objectives. We generally find that increasing the difficulty of the classification problem increases the overall quality of the embedding model: To that end, we introduce a novel generalization of mixup (Zhang et al., 2018) that mixes more than two sources (Section 2.1). Additionally, recent BirdCLEF competitions have demonstrated strong results with iterative pseudo-labeling (Kahl et al., 2024). Building on this observation, we introduce a variation of self-distillation (Allen-Zhu and Li, 2022) where a prototype learning classifier (Chen et al., 2019; Heinrich et al., 2025) is used to generate predictions that are used as soft targets for the network’s dense linear classifier (Section 2.2 and Section 2.3). Finally, noting that label granularity contributes to transfer learning performance (Section C.1), we add an auxiliary self-supervised loss in the form of source prediction (DIET, Balestriero, 2023), asking the model to predict the source recording of an audio window (Section 2.2 and Section 2.3). A similar objective has been shown to be useful for individual animal identification (Lapp et al., 2025).
@@ 2.5 Evaluation
The field of bioacoustics recognized domain shift as an important problem early on (Lasseck and Others, 2013), acknowledging that deployment conditions and tasks might differ from training time. Hence, our model selection (validation) is set up to test several forms of generalization: we consider avian soundscapes (which are qualitatively different from the focal recordings found in our training data); consider tasks other than species identification (e.g., call-type and dialect recognition); and evaluate transfer to species classification of non-avian taxa (bats, marine mammals, mosquitoes, etc.) For our model evaluation (testing) we use both BirdSet (Rauch et al., 2025b) and BEANS (Hagiwara et al., 2023). Combined, these benchmarks cover a similar set of domain shifts while still allowing us to make a fair comparison with existing models.
In practice, bioacoustics models like Perch are often used by practitioners who have limited computational resources and machine learning expertise while working on novel problems with little or no labeled data (but potentially large amounts of unlabeled data). In these settings a sensible approach is to embed the data once and then use methods such as clustering, nearest neighbor searches, few-shot learning and agile modeling (Dumoulin et al., 2025). We want to ensure that our model produces embeddings that perform well in these situations, so during both model selection and evaluation we freeze the embedding model.
@@ 2.5.1 Model-selection tasks
    We evaluate pretrained classifier performance using ROC-AUC on two fully-annotated bird datasets, where the class labels are a subset of the training classes of the Perch model: Powdermill (Denton et al., 2022) (as in Rauch et al. (2025b)) and Caples (Denton et al., 2022). These tasks validate that our model makes useful species predictions out-of-the-box. We make predictions for every 5 s window with a 2.5 s stride and count the prediction as correct when it overlaps with a ground truth annotation.
@@
    Finally, we have linear transfer tasks in which we select a random subset of $16$ examples per class and compute an embedding for each (by embedding each 5 s window with a stride of 5 s and averaging them). We then train a linear classifier for 10,000 steps on the selected examples using scikit-learn. All remaining examples are used to compute an ROC-AUC score. This follows the procedure applied in Ghani et al. (2023) and evaluates how well our model’s embeddings do in few-shot learning and agile modeling setups. We evaluate this task on the BEANS classification tasks, some of the datasets from Ghani et al. (2023) as well as marine datasets such as DCLDE (Palmer et al., 2025), NOAA (NOAA Pacific Islands Fisheries Science Center, 2021) and ReefSet (Williams et al., 2025).
@@
Powdermill Chronister et al. (2021); Rauch et al. (2025b)
@@
BirdSet Rauch et al. (2025b); Hopping et al. (2022); Vega-Hidalgo et al. (2023); Kahl et al. (2022a); Clapp et al. (2023); Kahl et al. (2022b)
@@ 2.5.2 Evaluation tasks
The BirdSet benchmark (Rauch et al., 2025b) consists of six fully-annotated datasets from the continental United States, Hawai‘i, Peru and Colombia. BirdSet provides training sets for each of these soundscapes for model fine-tuning. In our evaluations we forgo fine-tuning and simply use the predictions of our prototype learning classifier predictions directly.